# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v3

@@ 4.4.2. AI Impact on Humans and Society: how are humans influenced by AI systems ?
4.4.3. Evaluation in Human Studies: how might we evaluate and understand the impact of AI on humans and society?
We summarize common empirical methods used to rigorously understand and assess the impact of AI on humans. Specifically, we focus on two types of impact. On the micro-level, we discuss how to evaluate the effectiveness of human-AI collaboration; on the macro-level, we discuss how to assess the impact of AI on a large group of people over a long period of time.
Evaluate Human-AI Collaboration refers to the evaluation of the effectiveness of an AI system in collaboration with humans. It is key to not only consider the final output, but also the interaction experience (Lee et al., 2023b). Human-AI Team Performance compares human-AI team performance with the performance of humans alone without AI collaboration The metrics for measuring performance should include both task success metrics (e.g., accuracy) as well as indicators of efficiency (Mozannar et al., 2024; Ding et al., 2023a). Cognitive Workload and User Satisfaction involves understanding the degree of cognitive load that the user experiences when interacting with the system, as well as user satisfaction with the interaction and the final outcome. Such aspects are often captured via surveys or interviews (Ding et al., 2023a; Lee et al., 2023b). Control and Trust refers to how user control can support the avoidance of catastrophic AI failures, especially in high-stakes settings where AI mistakes could lead to harm (Srikanth et al., 2024; Ding et al., 2023b). Interaction Behavioral Analytics refers to measuring task performance quantitatively. This approach includes recording user interaction data and analyzing the patterns (Villalba et al., 2023; Perry et al., 2023). Qualitative Survey and Interviews refers to qualitative approaches to understanding human-AI interaction. Commonly used methods include qualitative survey questions (i.e., open-ended) and user interviews to assess aspects of the user experience. (Zhang et al., 2023b). Statistical Analysis utilizes methods such as regression analysis to quantitatively analyze and evaluate data from human studies, allowing for the verification of hypotheses (Wang et al., 2021).
Evaluate Societal Impact refers to the macro-impact of a group of people as they come to use AI broadly. This dimension requires sufficient scale and time. The aim is to understand how the group’s behavior changes as people within it frequently interact with AI. Public Opinion Surveys aims to investigate the impact of AI on human measures of interest though deploying and analyzing large scale questionnaires  (Santurkar et al., 2023). Behavioral Data Analytics collects large-scale and potentially longitudinal behavior data, with the aim of understanding how pattern evolve over time (Shin et al., 2023; Zhou and Lee, 2024).
    Underexplored Dimensions in Aligning Humans to AI (See Sec 5.1 for supporting data and evidence): 1. Educating and training humans on AI literacy is under-investigated. 2. Auditing AI for various ethical values is not fully explored. 3. The collaboration between humans and AI with similar or superior capabilities is under-explored. 4. The societal impacts of and reactions to AI advancements are not fully explored.
5. Findings and Discussions on Current Gaps
In this section, our aim is to consolidate key findings derived from our analysis of the framework and the current state of literature we reviewed. We begin by analyzing the overall trends and gaps in the literature (Section 5.1). We then focus on three essential aspects: the relationship between human values and alignment (Section 5.2), the potential interaction modes used to specify human values (Section 5.3).
@@ 5. Findings and Discussions on Current Gaps
Figure 4. The number of papers for each dimension in the bidirectional human-AI alignment framework. Out of papers that are relevant to each research question (i.e., gray bars), we show the number of papers that are relevant to each dimension (i.e., color bars). This figure illustrates the extent to which each dimension has been explored by existing research. We provide more analysis in Section 5.1.
5.1. Meta Analysis of Trends and Gaps
Based on our coding of all papers, we computed the number of relevant papers for each dimension in the proposed bi-directional framework. Based on the distribution in Figure 4, we noticed that in existing literature certain dimensions are over- or under-represented. We outline the under-explored dimensions in two directions in the highlighted grey bars in the end of Section 4.2 and Section 4.4, respectively, and provide more details below.
Underexplored dimensions in AI-centered alignment research. Most literature specified human values using explicit human feedback, whereas implicit and simulated human feedback were largely under-explored. In the integrating human specifications to AI research question, Besides, developing and customizing AI models during the inference stage or using the interactive approach is under-explored. Also, human-in-the-loop evaluation is much less explored than automatic evaluation.
Underexplored dimensions in Human-centered alignment research. Additionally, many studies on human cognitive adjustment to AI focus on enabling AI sensemaking and explanations so that humans can better understand, trust, and rely on AI. However, these studies often focused on explaining AI decision-making justification rather than educating people to acquire general skills and competencies required to understand, use, critique, and interact effectively with AI systems (i.e., AI literacy). AI literacy (Long and Magerko, 2020) plays a fundamental role in ensuring people understand and use AI correctly, and warrants deeper exploration. Also, we observed a wide range of studies (Wu et al., 2022; Gebreegziabher et al., 2023; Park et al., 2023; Ma et al., 2024) which developed interactive mechanisms and prototypes to empower human collaboration with AI. However, most studies assumed that AI plays an assistant role, being less capable than humans. This situation might change over time. Moreover, the influence of AI advancements on human behavior, social relationships, and societal changes is essential but remains largely unexplored.
Figure 5. The interaction techniques for specifying values in human-AI alignment. We compare the common interaction techniques used for the model “Learning” and “Inference” stages in human-focused (e.g., HCI) and AI-focused (e.g., NLP/ML) research studies.
    Takeaways of Interaction Techniques for Alignment. 1. Some common human feedback formats (rating, ranking) used in NLP/ML are not often studied in HCI. 2. Diverse human interactive feedbacks in HCI are not fully used in AI development in NLP/ML fields.
5.2. Insights into Human Values for Alignment
[…]
Universal vs. Personalized AI Values. While certain values are universally expected from AI (e.g., capability, equity, responsibility), others may be undesirable in specific contexts (Ngo et al., 2024) (e.g., seeking power). Simultaneously, AI models should be adaptable to diverse human value systems (Sorensen et al., 2024). Research is needed to develop methods for identifying appropriate value sets for specific individuals or groups, and for customizing AI to align with user values while maintaining ethical principles.
Disparities in Value Expectations and Evaluation. The fundamental differences between humans and AI necessitate distinct approaches to value evaluation. For instance, assessing AI honesty may require mechanistic interpretability (Bereska and Gavves, 2024), a more rigorous standard than that applied to humans. Future studies should explore methods for evaluating and explaining AI values and calibrating human expectations accordingly.
5.3. Interaction Techniques for Specifying Human Values: A Cross-Domain Analysis
Domain-Specific Interaction Techniques. The interaction techniques in AI-centered (NLP/ML) and Human-centered (HCI) alignment studies are often differ (Brath, 2021). NLP/ML studies primarily utilize numeric and natural language-based techniques. Also, NLP/ML research explore implicit feedback to extract human hidden feedback. In contrast, HCI research encompasses a broader range of graphical and multi-modal interaction signals (e.g., sketches, location information) beyond text and images. This disparity suggests potential gaps in extracting comprehensive human behavioral information.
6. Future Directions
@@
Drawing upon insights gained from the development of our framework and the associated coding analysis, as shown in Figure 6, we propose future research aiming to achieve the long-term alignment goal by identifying three important challenges from near-term to long-term objectives, including the Specification Game (Section 6.1), Dynamic Co-evolution of Alignment (Section 6.2), and Safeguarding Coadaptation (Section 6.3).
6.1. Specification Game
@@
[…]
Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors (Schwartz, 1994, 2012). Also, these priorities can change dynamically throughout an individual’s life stages. As such, It is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world (Gabriel, 2020). Future research, inspired by Social Choice Theory (Arrow, 2012), could focus on using democratic processes to aggregate individual values into collective agreements. Building on the summaries in Sections  4.1 and  5.2, researchers can employ democratic methods to identify diverse subsets of human values for AI alignment. Additionally, creating datasets that represent these values is crucial. Besides, it is crucial yet challenging for AI designers to investigate how to fully specify the appropriate values and to further integrate these values into AI alignment. Future important area involves developing algorithms, such as the Bradley-Terry Model (Rafailov et al., 2023) or Elo Rating System (Bai et al., 2022b), to convert heterogeneous human values into AI-compatible formats for training reward models and guiding reinforcement learning. Researchers should also explore AI models capable of aligning with unstructured human data, including free-form descriptions of values, multimedia, or sensor recordings depicting human behavior.
[…]
Current alignment methods use instructions, ratings, and rankings to infer human values, which can not fully capture all relevant human values and constraints. Future research should focus on optimizing interactive interfaces to efficiently elicit human values. These interfaces can leverage diverse interaction modes to capture comprehensive human value information. Additionally, people often struggle to formulate optimal prompts for AI, accurately specify their requirements, and articulate their desired values, which can change based on context and time. Developing proactive interfaces that use conversational techniques to elicit nuanced and evolving values is also crucial. Implicit human signals that indicate values are also frequently overlooked. Additionally, systems that track interactions to hypothesize and validate implicit human values in real-time should be designed.
6.2. Dynamic Co-evolution of Alignment

### v4

@@ 3 Bidirectional Human-AI Alignment Framework
Evaluation in Human Studies how might we evaluate the impact of AI on humans and society? Evaluate Human-AI Collaboration Lee2022EvaluatingHM ; ding-etal-2023-harnessing
Evaluate Societal Impact santurkar2023whose ; Shin2023SuperhumanAI
This section introduces the Bidirectional Human-AI Alignment framework, which encompasses two interconnected alignment directions as a feedback loop, as shown in Figure 1. The “Align AI to Humans” direction refers to mechanisms to ensure that AI systems’ values match those of humans’. The “Align Humans to AI” direction investigates human cognitive and behavioral adaptation to AI advancement. We introduce more details below.
3.1 Align AI to Humans
This direction delineates alignment research from AI-centered perspective (e.g., ML/NLP domains) and provides AI developers and researchers with approaches for handling two main challenges: carefully specifying the values of the system, and ensuring that the system adopts the specification robustly wiki_AI_alignment ; ngo2022alignment . Therefore, as shown in Table 1, we explore two core research questions in this direction as: RQ1. What relevant human values are studied for AI alignment, and how do humans specify these values? and RQ2. How can human values be integrated into the AI systems?
  • RQ1: Human Values and Specifications.
To identify key human values and specifications for AI alignment, we begin by addressing two critical subquestions: (1) what values have AI systems been aligned with? and (2) how can humans interactively specify values during AI development? As summarized at the top of Table 1, we present the key dimensions that emerged from our analysis of these questions, along with supporting studies.
Categorizing Aligned Human Values. To systematically understand human values relevant to human-AI alignment shen2024valuecompass ; shen2025mind , we draw on the adapted Schwartz Theory of Basic Values, examining values along two dimensions: Sources and Types. The Sources dimension categorizes values as individual (e.g., personal interests and biological needs like factuality or cognitive biases) schwartz1994there ; kiesel2022identifying ; jin2022make , social (e.g., shared group norms such as fairness or morality) schwartz1994there ; ammanabrolu2022aligning ; santy2023nlpositionality , and interactive (e.g., expectations in human-AI interactions like usability, autonomy, and trust) hovy2021importance ; cabrera2023zeno ; wang2023humanoid . The Types dimension organizes values into four high-order categories: Self-Enhancement (e.g., achievement, power) sedikides1995multiply ; he2023interaction ; xu2023lemur , Self-Transcendence (e.g., benevolence, honesty, fairness) frankl1966self ; banovic2023being ; wu2023cross , Conservation (e.g., safety, tradition, conformity) conservatism ; qi2023fine ; zagalsky2021design , Openness to Change (e.g., creativity, privacy, autonomy) bakker2022fine ; lima2020collecting ; ashktorab2020human . These dimensions offer a comprehensive framework for evaluating and aligning AI systems with the multifaceted nature of human values.
Interaction Techniques to Specify AI Values. This sub-research question explores how human values are interactively specified to ensure AI alignment, focusing on the techniques through which AI systems manifest or internalize these values. It identifies three main approaches: explicit human feedback, where values are directly communicated via principles, ratings, natural language interactions, or multimodal inputs like gestures and images petridis2024constitutionmaker ; dai2023safe ; bai2022constitutional ; implicit human feedback, where values are inferred from indirect cues such as discarded options, language patterns, theory of mind reasoning, and social relationships isajanyan2024social ; park2020scalable ; lu2022towards ; and simulated human value feedback, where AI systems approximate human responses using feedback simulators, comparisons to human data, or synthetically generated data petridis2024constitutionalexperts ; dong2023aligndiff ; kim2023aligning . Together, these approaches illuminate the mechanisms by which AI systems interpret and enact human values via direct and indirect human-AI interaction.
Key Takeaways. By comparing prior research with our comprehensive analysis of human values and interaction techniques, we found that existing studies are largely constrained to conventional principles and standard interaction methods, overlooking the broader spectrum of human values and the interactive approaches needed to specify them effectively in alignment.
  • RQ2: Integrating Human Specifications into AI.
Building on the value-laden human specifications gathered through interaction, we next explore diverse methods for integrating human values into AI systems. We examine this central question across two key stages of the AI lifecycle—development and deployment—by asking: (1) how can general or customized human values be integrated throughout the AI development process? and (2) what methods and platforms are available to evaluate values during AI development?
Integrating General Values to AI. This sub-research question examines how broad, universally recognized human values are embedded into AI systems to ensure ethical alignment and societal acceptance. It outlines three key dimensions: Instruction Data, which includes human annotations, human-AI co-annotation, and simulated human data to guide value-based training shen2023multiturncleanup ; sun2024principle ; dubois2024alpacafarm ; Model Learning, where human values are integrated during training through either real-time online alignment or offline processes prior to deployment ouyang2022training ; dong2023raft ; rafailov2024direct ; and Inference Stage, where AI systems are evaluated and refined using techniques such as prompting, external tool interactions, and response search to ensure their outputs align with predefined ethical criteria bai2022constitutional ; wang2023humanoid ; gou2023critic ; bradley2023quality . Together, these processes aim to build AI systems that promote trust and responsible use.
Customizing AI Values. This sub-research question investigates how AI systems can be customized to reflect specific user preferences, application domains, or community values, thereby improving contextual alignment. It identifies three primary strategies: Customized Data, which involves curating and finetuning datasets based on socio-demographic groups, user histories, or expert selections to align models with targeted human values felkner2023winoqueer ; orlikowski2023ecological ; lyu2023exploiting ; Adapt Model by Learning, which includes techniques such as group-based learning, active learning, adapter insertion, mixture of experts, and enhanced knowledge integration to refine model behavior zhao2023group ; peschl2022moral ; sun2022bertscore ; and Interactive Alignment, which actively engages users through real-time feedback, steering prompts, and proactive adjustments based on user profiles to tailor AI systems to specific contexts and preferences maghakian2022personalized ; swazinna2022user ; lahoti2023improving .
Evaluating AI Systems. This sub-research question examines how the integration of human values into AI systems, particularly large language models (LLMs), is evaluated, highlighting both human-in-the-loop and automated methods. Human-in-the-loop evaluation involves human judgment, feedback, and collaboration to assess the ethical and value alignment of AI outputs, leveraging direct human input or combined human-AI assessment processes sharma2023cognitive ; ouyang2023shifted ; ammanabrolu2022aligning . Automatic evaluation utilizes computational techniques, including human simulators, standardized benchmarks, and distributional comparisons, to evaluate alignment without human intervention jia2024towards ; sun2024principle ; patel2021stated . Together, these approaches aim to ensure that AI systems reflect human ethical standards and value frameworks effectively and reliably.
Ecosystem and Platforms. The ecosystem and platforms refer to the broader context in which AI systems operate and interact with other agents, platforms, or environments. This includes the infrastructure, frameworks, and technologies that support the development, deployment, and utilization of AI systems. LLM-based Agents are based on large language models (LLMs) such as GPT (Generative Pre-trained Transformer) models, which have been pre-trained on vast amounts of text data dubois2024alpacafarm ; xu2023gentopia ; wu2023autogen ; zhou2023sotopia . RL-based Agents are based on reinforcement learning (RL) algorithms to learn and adapt their behavior based on feedback from the environment or human users yuan2024uni . Annotation Platforms refers to the ecosystems that are designed to crowdsource human demonstrations as collected data for reinforcement learning gerstgrasser2021crowdplay and supervised finetuning learning for alignment pei2022potato .
Key Takeaways. Our systematic analysis of value integration and evaluation in AI reveals a strong focus on explicit value annotations, while implicit value expressions and behaviors are often neglected. Despite the power of general-purpose AI, methods for customizing systems to reflect individual or group values remain underexplored. Additionally, there is a lack of standardized criteria for evaluating human-in-the-loop methods and supporting platforms, underscoring the need for more robust and context-sensitive evaluation frameworks in value-aligned AI development.
3.2 Align Humans to AI
From a long-term perspective, it is crucial to consider the dynamic and evolving nature of human-AI alignment. This direction emphasizes a human-centered perspective—drawing from fields such as HCI and the social sciences—and offers guidance for researchers and user experience designers in addressing two core research questions: RQ3. How can humans learn to perceive, explain, and critique AI? and RQ4. How do individuals and society adapt their behaviors in response to AI advancements?
  • RQ3: Human Cognitive Adjustment to AI.
For effective collaboration and value specification, humans need to develop a clear understanding of how AI systems function. As AI introduces various risks, fostering critical thinking is essential to prevent blind reliance. To address this, we systematically investigate, as summarized in Table 1, (1) how humans learn to perceive and understand AI, and (2) how they can engage in critical reflection on AI behavior and outputs.
Perceiving and Understanding AI. This sub-research question explores how to enhance human understanding and perception of AI systems, particularly among non-technical users, through education, training, and human-centered explanation techniques. It emphasizes the importance of AI literacy and awareness as foundational competencies for effective human-AI collaboration long2020ai , supported by explicit training courses designed to improve users’ ability to engage with AI ketan-ai-medicine-education ; mcdonald2020intersectional . Additionally, it highlights efforts in AI sensemaking and human-centered explanations, including visualizations and interactive techniques, to help people interpret AI mechanisms and outputs, thereby fostering more informed and meaningful human-AI interactions kaur2022sensible ; sivaraman2023ignore ; tenney2020language .
Critical Thinking around AI. This sub-research question examines how individuals critically reflect on and evaluate AI systems by comparing their own mental models with those of the AI, focusing on the rationality, reliability, and ethical behavior of these technologies. It emphasizes the need for humans to develop critical thinking skills to identify biases, errors, and ethical concerns in AI outputs, and to audit AI systems for compliance with moral and societal standards. Key areas include building appropriate levels of Trust and Reliance on AI based on its competence and reliability yin2019understanding ; ma2023should ; schemmer2023appropriate ; clark2025epistemic , engaging in selective AI adoption aligned with user needs and values gu2024ai ; vasconcelos2023explanations , and addressing Ethical Considerations through mitigation strategies and auditing holstein2019improving ; bandy2021problematic . Additionally, it underscores the importance of Recalibrating Cognition, where users adjust their understanding and expectations of AI performance and reliability to foster a balanced and informed relationship with AI systems he2023interaction ; kocielnik2019will ; wischnewski2023measuring .
Key Takeaways. Our systematic review reveals a significant gap in research on training and educating humans to develop appropriate knowledge, trust, and reliance when collaborating with AI systems. Additionally, there is limited exploration of how to critically evaluate AI across a broad spectrum of human values.
  • RQ4: Human Adaptation Behavior to AI.
Building on our investigation of human cognitive adjustments to AI, we further examine how individuals and society can respond effectively to AI’s expanding influence. To guide this inquiry, we address three key questions: (1) How do humans learn to collaborate with AI across its diverse roles? (2) In what ways are individuals and society affected by AI? and (3) How can these impacts be comprehensively assessed? The key dimensions emerging from this analysis are summarized at the bottom of Table 1.
Human-AI Collaboration Mechanisms. This category explores the diverse ways humans and AI collaborate through partnerships, co-creation, and mutual learning. AI Assistants, particularly those powered by LLMs, support human tasks by interpreting user demands, enhancing prompt formulation, generating creative prototypes, and aiding decision-making wu2022ai ; gebreegziabher2023patat ; scattershot . In collaborative frameworks, humans and AI function as AI Partners, where simulated agency and reciprocal learning enable joint decision-making and knowledge sharing park2023generative ; zagalsky2021design ; bernstein2010soylent . AI delegation and mediation transform traditional human tasks, while co-design with AI treats the system as both a collaborator and a design material. Additionally, AI Tutoring systems enhance human learning in both technical and social domains by offering tailored feedback, adaptive instruction, and immersive practice environments, ultimately improving skill acquisition and performance ma2023hypocompass ; peng2021say ; shaikh2023rehearsal .
AI Impact on Humans and Society. This category investigates the multifaceted effects of AI advancement on human behavior, attitudes, and societal dynamics, aiming to inform policy, education, and interventions. The Impacts on Participatory Individuals and Groups dimension focuses on how AI influences human decision-making, creativity, privacy, and authorship—shaping behaviors and raising concerns about data rights and intellectual ownership shen2020useful ; ashkinaze2024ai ; li2024human ; shi2025siren . The Societal Concerns and AI Impacts dimension expands this lens to the societal level, examining AI’s effects on misinformation, education, social norms, and the workplace, including issues like disinformation, shifts in learning practices, changes in interpersonal relationships, and job displacement 10.1145/3610082 ; kazemitabaar2023novices ; cheng2022human . The Reaction to AI Advancement dimension explores regulatory, cultural, and institutional responses to AI, addressing how societies perceive, govern, and adapt to AI technologies lima2023blaming ; hacker2023regulating ; 10.1145/3593013.3593992 . This includes efforts to regulate bias and discrimination, develop policy frameworks, track evolving AI acceptance, ensure transparency and oversight, and establish responsible AI checklists to guide ethical and safe deployment.
Evaluation in Human Studies. This summary covers common empirical methods used to rigorously evaluate AI’s impact on humans at both micro and macro levels. At the micro-level, Human-AI Collaboration Evaluation assesses not only task success and efficiency but also user experience, including cognitive workload, user satisfaction, control, and trust, especially in critical settings to prevent failures Lee2022EvaluatingHM ; Mozannar2024TheRE ; ding-etal-2023-harnessing ; Lee2022EvaluatingHM . Methods include quantitative interaction analytics, qualitative surveys and interviews, and statistical analyses to understand and verify user behaviors and perceptions. At the macro-level, Societal Impact Evaluation focuses on understanding long-term behavioral changes within large populations as AI use becomes widespread, employing large-scale public opinion surveys and behavioral data analytics over time to capture evolving patterns and societal shifts related to AI interaction santurkar2023whose ; Shin2023SuperhumanAI ; Zhou2024GenerativeAI .
Key Takeaways. Our review reveals that while prior studies have extensively examined how AI can assist humans across various tasks, there is limited exploration of the challenges humans face when collaborating with AI that surpass human capabilities in certain domains. Additionally, more research is needed to understand the evolving impact of AI on individuals and society at large over time.
4 Underexplored Research Gaps and Challenges
@@
Figure 2: The number of papers for each dimension in the bidirectional human-AI alignment framework. Out of papers that are relevant to each research question (i.e., gray bars), we show the number of papers that are relevant to each dimension (i.e., color bars).
In this section, we consolidate key findings from our systematic analysis of the framework and the reviewed literature (in Figure 2. To accurately capture the distribution of existing research and identify current gaps and challenges, we quantified the number of relevant studies corresponding to each dimension in the Bidirectional Human-AI Alignment. We elaborate on key findings below.
Underexplored Dimensions in Aligning AI with Humans. Current AI alignment research has primarily focused on incorporating explicitly stated human values, often gathered through direct feedback mechanisms such as ratings, rankings, or instructions. However, several critical dimensions remain underexplored. First, the use of implicit human feedback — such as behavioral cues, physiological signals, or interaction patterns—and simulated human value feedback has received limited attention, despite their potential to provide rich, context-sensitive information about human values. Second, most alignment efforts concentrate on model training stages, whereas developing and customizing AI models during inference or through interactive processes for embedding and evaluation values remains significantly underexplored. Enabling real-time adaptation of AI behavior to human input could enhance alignment in dynamic or personalized contexts. Third, human-in-the-loop evaluation, which involves assessing AI systems through active human participation and feedback, is rarely used in comparison to fully automated evaluation metrics. Expanding research into these areas is essential for advancing more robust, responsive, and context-aware alignment.
Underexplored Dimensions in Aligning Humans with AI. Human-centered alignment research has predominantly emphasized designing AI systems that facilitate human understanding through sensemaking and explanation—primarily by clarifying the justifications behind AI decisions to foster user trust and reliance. However, this focus often overlooks the broader goal of fostering AI literacy—the essential skills and competencies individuals need to understand, critique, use, and interact effectively with AI systems. Despite its foundational role in responsible AI engagement, AI literacy remains an underexplored area. Furthermore, while numerous studies have proposed interactive mechanisms and prototypes to support human-AI collaboration, they commonly assume that AI operates in a subordinate or assistive role. As AI systems grow increasingly capable, research must also consider collaborative dynamics between humans and AI with equal or superior capabilities. Additionally, the ethical auditing of AI from a human-centered perspective and the societal-level impacts of AI — such as changes in human behavior, social relationships, and public responses—have not been sufficiently examined. These dimensions are critical for ensuring meaningful and equitable alignment between humans and evolving AI technologies.
5 Near to Long-term Risks and Opportunities
@@
Drawing upon insights gained from the development of our framework and the associated systematic review analysis, we propose future research aiming to achieve the long-term alignment goal by identifying three important challenges from near-term to long-term objectives, including the Specification Game, Dynamic Co-evolution of Alignment, and Safeguarding Coadaptation.
5.1 Specification Game
@@
Integrate fully specified human values into aligning AI. Individuals often possess value systems that encompass multiple values with varying priorities, rather than a single value, to guide their behaviors schwartz1994there ; schwartz2012overview . Also, these priorities can change dynamically throughout an individual’s life stages. As such, It is more realistic to select values compatible with specific societies or situations, given the fact that we live in a diverse world gabriel2020artificial . Future research, inspired by Social Choice Theory arrow2012social , could focus on using democratic processes to aggregate individual values into collective agreements. Building on the summaries in Sections  3.1, researchers can employ democratic methods to identify diverse subsets of human values for AI alignment. Additionally, creating datasets that represent these values is crucial. Besides, it is crucial yet challenging for AI designers to investigate how to fully specify the appropriate values and to further integrate these values into AI alignment. Future important area involves developing algorithms, such as the Bradley-Terry Model rafailov2024direct or Elo Rating System bai2022constitutional , to convert heterogeneous human values into AI-compatible formats for training reward models and guiding reinforcement learning. Researchers should also explore AI models capable of aligning with unstructured human data, including free-form descriptions of values, multimedia, or sensor recordings depicting human behavior.
Elicit nuanced and contextual human values during diverse interactions. Current alignment methods use instructions, ratings, and rankings to infer human values, which can not fully capture all relevant human values and constraints. Future research should focus on optimizing interactive interfaces to efficiently elicit human values. These interfaces can leverage diverse interaction modes to capture comprehensive human value information. Additionally, people often struggle to formulate optimal prompts for AI, accurately specify their requirements, and articulate their desired values, which can change based on context and time. Developing proactive interfaces that use conversational techniques to elicit nuanced and evolving values is also crucial. Implicit human signals that indicate values are also frequently overlooked. Additionally, systems that track interactions to hypothesize and validate implicit human values in real-time should be designed.
5.2 Dynamic Co-evolution of Alignment